OpenStudy (anonymous):

Find all solutions to the equation. (sin x)(cos x) = 0

OpenStudy (anonymous):

@zepdrix @Cha1234

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(sin x)(cos x) = 0 sin(x) =0 or cos(x) = 0 x = 0, pi, 2pi, .... or x = pi/2, 3pi/2, ... x = pi*n or x = pi/2 + pi*n combine the two solution sets to get x = n*pi/2 Note: the value of n is an integer